OpenStudy (anonymous):

please help The length of a rectangle is 8 in. more than its width. The perimeter is 52 in. What is the length of the rectangle?

OpenStudy (anonymous):

The perimeter of the rectange is equal to: P=2*(width + length)=2(W + L) We know that length is 8 in more than width, so L=W + 8 in. We also know that P=52in. So here is our equation: 52=2(W + W + 8) Now simplify

OpenStudy (anonymous):

are we figuring out W? if so w = 9

OpenStudy (anonymous):

Combine like terms in brackets: 52=2(2W+8) Divide both sides by 2: 52/2 = 2/2(2W+8) Simplify: 26=2W+8 Subtract 8 from both sides: 26-8=2W+8-8 Simplify: 18=2W Divide both sides by 2: 18/2=2W/2 Simplify: 9=W We got that width is 9 inches. So now solve for length: L=W+8 L=9+8=17 Inches The Answer is: The length of the rectangle is 17 inches

OpenStudy (anonymous):

We could figure out L directly, by wrting W=L-8 inches instead of L=W+8 inches and so we will have another equation, but the answer will be the same

OpenStudy (anonymous):

thank you.. im new to openstudy but i would like to medal. how?

OpenStudy (anonymous):

click on bet response near my message and you are done! :)

OpenStudy (anonymous):

*best response